Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 6.8-in and a standard deviation of 0.8-in. Due to financial constraints, the helmets will be designed to fit all men except those with head breadths that are in the smallest 0.7% or largest 0.7%.

What is the minimum head breadth that will fit the clientele?
min =

What is the maximum head breadth that will fit the clientele?
min =

Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

z score corresponding to 0.7% area to its left = -2.457

z score corresponding to 0.7% area to its right = 2.457

Hence,

Minimum head breadth = 6.8 - 2.457*0.8 = 4.8

Maximum head breadth = 6.8 + 2.457*0.8 = 8.8